# Closed-Form Approximation of the Total Variation Proximal Operator

### v1

@@ Abstract
Total variation (TV) is a widely used function for regularizing imaging inverse problems that is particularly appropriate for images whose underlying structure is piecewise constant. TV regularized optimization problems are typically solved using proximal methods, but the way in which they are applied is constrained by the absence of a closed-form expression for the proximal operator of the TV function. A closed-form approximation of the TV proximal operator has previously been proposed, but its accuracy was not theoretically explored in detail. We address this gap by making several new theoretical contributions, proving that the approximation leads to a proximal operator of some convex function, that it always decreases the TV function, and that its error can be fully characterized and controlled with its scaling parameter. We experimentally validate our theoretical results on image denoising and sparse-view computed tomography (CT) image reconstruction.
@@ 2 Related Work
The TV regularizer promotes sparse image gradients, producing reconstructed images that are approximately piecewise constant. TV regularization has been demonstrated to be effective in various inverse problems, including image deconvolution, diffraction ultrasound tomography, compressed sensing, and optical tomography [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. APGM [3] is often employed to solve image reconstruction problems with the TV regularizer [5]. However, APGM requires additional sub-iterations to compute the proximal operator of the TV regularizer, which can slow down the overall convergence. While ADMM with an appropriate variable-splitting strategy can be used to avoid these sub-iterations [13], the introduction of an additional auxiliary variable representing the image gradients increases memory requirements.
@@ 3.1 Method
${\bm{W}}:=\frac{1}{2\sqrt{d}}\begin{bmatrix}{\bm{A}}\vspace{3pt}\\ {\bm{D}}\end{bmatrix},\qquad\begin{aligned} {\bm{A}}&:=\big{[}{\bm{A}}_{1}^{{\mathsf{T}}}~{}\cdots~{}{\bm{A}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\\ {\bm{D}}&:=\big{[}{\bm{D}}_{1}^{{\mathsf{T}}}~{}\cdots~{}{\bm{D}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\end{aligned}$
where ${\bm{A}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ represents an averaging operator along the $j^{\text{th}}$ dimension, and ${\bm{D}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ denotes a discrete gradient operator along the $j^{\text{th}}$ dimension. ${\bm{A}}_{j}$ and ${\bm{D}}_{j}$ are convolutions along the $j^{\text{th}}$ dimension of the signal with the kernels
${\bm{a}}=\begin{bmatrix}1\vspace{3pt}\\ 1\end{bmatrix},\qquad{\bm{d}}=\begin{bmatrix}1\vspace{3pt}\\ -1\end{bmatrix},$
respectively. The mapping ${\bm{W}}$ can be interpreted as a union of scaled and shifted first-level discrete Haar wavelet and scaling functions across all $d$ dimensions [9]. It is important to note that this operator satisfies ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}=\bm{I}$, but due to redundancy in the union of scaled orthogonal transforms, ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}\neq\bm{I}$ [54].
@@
${\mathcal{S}}_{\tau}({\bm{z}})\,\coloneqq\,{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.04996pt2\sqrt{d}}({\bm{W}}{\bm{z}}),$ (6)
where ${\mathcal{T}}_{\lambda}$ is an operator with the parameter $\lambda\,\coloneqq\,\tau\hskip 1.49994pt2\sqrt{d}>0$, corresponding to the application of component-wise shrinkage functions on the scaled difference components of ${\bm{W}}{\bm{z}}$. The anisotropic shrinkage function ${\mathcal{T}}^{\hskip 1.04996pta}_{\lambda}$ is defined as
${\mathcal{T}}^{\hskip 1.04996pta}_{\lambda}(u)\,\coloneqq\,{\mathsf{max}}(|u|-\lambda,0)\frac{u}{|u|},$ (7)
where $u:=1/\big{(}2\sqrt{d}\big{)}[{\bm{D}}{\bm{z}}]_{j}\in\mathbb{R}$ for $j\in\{1,2,...,nd\}$. The isotropic shrinkage function ${\mathcal{T}}^{\hskip 1.04996pti}_{\tau}$ is defined as
${\mathcal{T}}^{\hskip 1.04996pt\text{i}}_{\lambda}({\bm{u}}):={\mathsf{max}}(\|{\bm{u}}\|_{2}-\lambda,0)\frac{{\bm{u}}}{\|{\bm{u}}\|_{2}},$ (8)
with ${\bm{u}}:=1/\big{(}2\sqrt{d}\big{)}\big{[}[{\bm{D}}_{1}{\bm{z}}]_{i}~{}\cdots~{}[{\bm{D}}_{d}{\bm{z}}]_{i}\big{]}^{\mathsf{T}}\in\mathbb{R}^{d}$ for $i\in\{1,2,\cdots,n\}$. Note that the shrinkage parameter $\tau$ must be scaled by $2\sqrt{d}$ due to the scaling of ${\bm{W}}$. For simplicity, we omit the superscripts a and i when we do not need to distinguish between the anisotropic and isotropic thresholding functions, and the corresponding TV functions. The following results apply to both versions.
@@ 3.2 Theoretical Analysis
We will next demonstrate that $\widehat{h}$ provides a close approximation of ${h}$. To effectively describe the operator as an approximation of the true proximal operator for ${h}$, we introduce the concept of the subdifferential [55, 56].
@@
Proposition 2 (b) and 2 (c) rely on notions of approximate proximal operators, commonly used in the literature. Specifically, Proposition2 (b) aligns with a “type 3 approximation” in the proximal operator with $\tau\epsilon_{1}$-precision, while Proposition 2 (c) corresponds to a “type 2 approximation” with $\tau\sqrt{2\epsilon_{2}}$-precision [48].
Proposition 2 (a) indicates that the operator decreases the TV function, meaning ${\mathcal{S}}_{\tau}({\bm{z}})$ serves as a valid descent direction for TV. However, it does not necessarily provide the optimal descent direction that would be obtained from the true proximal operator of ${h}$. The second part of Proposition 2 demonstrates that the operator functions as a perturbed proximal operator of TV. Specifically, the output of the operator is equivalent to applying the true proximal operator of TV to a perturbed input vector ${\bm{z}}$, modified by a vector ${\bm{\delta}}$. Notably, the norm of the perturbation $\|{\bm{\delta}}\|_{2}$ is bounded and can be controlled by adjusting $\tau$. From the optimality condition for the proximal operator of true TV regularizer ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$, there is the following well known relationship
[…]
The third part of Proposition 2 establishes an analogous connection for the closed-form approximate operator, characterized using the $\epsilon$-subdifferential from Definition 2, where $\epsilon$ is controlled by the proximal scaling parameter $\tau$.
Overall, Proposition 2 implies that the $\widehat{h}$ from Proposition 1 can be seen as an approximation of the true TV function ${h}$ with controllable accuracy. Notably, by reducing $\tau$, one can achieve an arbitrarily close approximation to the true TV proximal operator. Therefore, our theoretical result supports the use of ${\mathcal{S}}_{\tau}({\bm{z}})$ as an approximation of ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$.
Corollary 1.
For any $\tau>0$ and for any ${\bm{z}}\in\mathbb{R}^{n}$
$\big{\|}{\mathsf{prox}}_{\tau{h}}({\bm{z}})-{\mathcal{S}}_{\tau}({\bm{z}})\big{\|}_{2}\leq\tau\epsilon_{1}\;.$
Corollary 1 states that the $\ell_{2}$ difference between the true TV proximal operator and the approximate operator ${\mathcal{S}}_{\tau}({\bm{z}})$ has a bounded distance. As the accuracy of the operator is controlled by $\tau$, it might initially appear that it would only be effective for problems with a very small regularization parameter $\lambda$ for TV. However, this is not the case when ${\mathcal{S}}_{\tau}({\bm{z}})$ is used within iterative algorithms such as APGM and ADMM. As shown in Algorithms 1 and 2, the shrinkage parameter for ${\mathcal{S}}_{\tau}({\bm{z}})$ is set by $\tau=\gamma\lambda$. Thus, one can enhance accuracy by adjusting $\gamma$ rather than altering the TV regularization parameter $\lambda$ For APGM, $\gamma$ represents the step size, while for ADMM, $\gamma$ serves as the penalty parameter in the augmented Lagrangian.
@@
Table 1: The relative cost, PSNR with respect to the exact TV solution and groundtruth for various regularization ($\lambda$) and step-size/penalty ($\gamma$) parameters for image denoising and CT reconstruction. For the APGM and ADMM algorithms, $\tau=\gamma\lambda$, where $\gamma$ determines the accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$. The metrics have been averaged across the 10 images in the dataset. Note that $L$ is the Lipschitz constant of $\nabla g$.
[…]
@@ 4 Numerical Validation
In this section, we numerically validate Proposition 2 by comparing the reconstruction performance of iterative proximal algorithms when ${\mathcal{S}}_{\tau}({\bm{z}})$ replaces the true proximal operator. Specifically, we run APGM and ADMM on a simulated low-angle computed tomography (CT) image reconstruction problem. The APGM and ADMM algorithms using ${\mathcal{S}}_{\tau}({\bm{z}})$ instead of the true proximal operator are shown in Algorithms 1 and 2, respectively. The accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$ can be directly controlled by $\gamma$ since $\tau=\gamma\lambda$, where $\gamma$ is the step size and penalty parameter, for APGM and ADMM algorithms, respectively. To calculate the true proximal operator, we use the iterative Fast Projected Gradient (FPG) algorithm [5]. Additionally, we investigate the performance of the closed-form approximate operator relative to true proximal operator of TV for an image denoising task. The numerical evaluations are performed using 10 piecewise-smooth foam phantoms generated by the XDesign Python package [58] and the implementationsSee classes AnisotropicTVNorm and IsotropicTVNorm. of ${\mathcal{S}}_{\tau}({\bm{z}})$ in the SCICO Python package [59]. All algorithms are run until the stopping criteria

### v2

@@ Abstract
Total variation (TV) is a widely used function for regularizing imaging inverse problems that is particularly appropriate for images whose underlying structure is piecewise constant. TV regularized optimization problems are typically solved using proximal methods, but the way in which they are applied is constrained by the absence of a closed-form expression for the proximal operator of the TV function. A closed-form approximation of the TV proximal operator has previously been proposed, but its accuracy was not theoretically explored in detail. We address this gap by making several new theoretical contributions, proving that the approximation leads to a proximal operator of some convex function, it is equivalent to a gradient descent step on a smoothed version of TV, and that its error can be fully characterized and controlled with its scaling parameter. We experimentally validate our theoretical results on image denoising and sparse-view computed tomography (CT) image reconstruction.
@@ 2 Related Work
The TV regularizer promotes sparse image gradients, producing reconstructed images that are approximately piecewise constant. TV regularization has been demonstrated to be effective in various inverse problems, including image deconvolution, diffraction ultrasound tomography, compressed sensing, and optical tomography [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. APGM [3] is often employed to solve image reconstruction problems with the TV regularizer [5]. However, APGM requires additional sub-iterations to compute the proximal operator of the TV regularizer, which can slow down the overall convergence. Other approaches such as Primal-Dual Hybrid Gradient Algorithm (PDHG) [30] and ADMM [13] can be used to avoid these sub-iterations; but they have their own drawbacks such as increased memory consumption to represent the image gradients.
@@ 3.1 Method
${\bm{W}}:=\frac{1}{2\sqrt{d}}\begin{bmatrix}{\bm{M}}\vskip 3.0pt\\ {\bm{D}}\end{bmatrix},\qquad\begin{aligned} {\bm{M}}&:=\big{[}{\bm{M}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{M}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\\ {\bm{D}}&:=\big{[}{\bm{D}}_{1}^{{\mathsf{T}}}~\cdots~{\bm{D}}_{d}^{{\mathsf{T}}}\big{]}^{\mathsf{T}},\end{aligned}$
where ${\bm{M}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ represents an averaging operator along the $j^{\text{th}}$ dimension, and ${\bm{D}}_{j}:\ \mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ for $j\in\{1,...,d\}$ denotes a discrete gradient operator along the $j^{\text{th}}$ dimension. ${\bm{M}}_{j}$ and ${\bm{D}}_{j}$ are convolutions along the $j^{\text{th}}$ dimension of the signal with the kernels
${\bm{m}}=\begin{bmatrix}1\vskip 3.0pt\\ 1\end{bmatrix},\qquad{\bm{d}}=\begin{bmatrix}1\vskip 3.0pt\\ -1\end{bmatrix}$,
respectively. In this work, we assume periodic boundary conditions for both $\bm{M}$} and $\bm{D}$}. Note that this is not essential for the method or the results in this paper: aperiodic boundary conditions are perfectly acceptable. The mapping $\bm{W}$} can be interpreted as a union of scaled and shifted first-level discrete Haar wavelet and scaling functions across all $d$ dimensions [9]. It is important to note that this operator satisfies ${\bm{W}}^{{\mathsf{T}}}{\bm{W}}=\bm{I}$, but due to redundancy in the union of scaled orthogonal transforms, ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}$ is over-determined and so ${\bm{W}}{\bm{W}}^{{\mathsf{T}}}\neq\bm{I}$ [55] (Section 3, first paragraph).
@@
${\mathcal{S}}_{\tau}({\bm{z}})\,\coloneqq\,{\bm{W}}^{{\mathsf{T}}}{\mathcal{T}}_{\tau\hskip 1.22911pt2\sqrt{d}}({\bm{W}}{\bm{z}}),$ (6)
where ${\mathcal{T}}_{\lambda}$ is an operator with the parameter $\lambda\,\coloneqq\,\tau\hskip 1.49994pt2\sqrt{d}>0$, corresponding to the application of component-wise shrinkage functions on the scaled difference components of ${\bm{W}}{\bm{z}}$. For the anisotropic operator ${\mathcal{T}}^{\hskip 1.22911pta}_{\tau 2\sqrt{d}}$, the shrinkage function
${\mathsf{max}}(|u|-\lambda,0)\frac{u}{|u|},$ (7)
is applied to each difference component; that is, $u=[{\bm{W}}{\bm{z}}]^{\text{dif}}_{i}\in\mathbb{R}$ for $i\in\{1,2,...,nd\}$. For the isotropic operator ${\mathcal{T}}^{\hskip 1.22911pti}_{\tau 2\sqrt{d}}$, the shrinkage function
${\mathsf{max}}(\|{\bm{u}}\|_{2}-\lambda,0)\frac{{\bm{u}}}{\|{\bm{u}}\|_{2}},$ (8)
is applied to each finite difference vector corresponding to each pixel in $\bm{z}$}; that is, ${\bm{u}}=[{\bm{W}}{\bm{z}}]^{\text{dif}}_{G_{i}}\in\mathbb{R}^{d}$ for $i\in\{1,2,\cdots,n\}$. Each $G_{i}$ contains the $d$ indices in ${\bm{W}}{\bm{z}}]^{\text{dif}}$ corresponding to the components of the finite difference vector at pixel $i$.
Note that the shrinkage parameter $\tau$ must be scaled by $2\sqrt{d}$ due to the scaling of $\bm{W}$}. For simplicity, we omit the superscripts a and i when we do not need to distinguish between the anisotropic and isotropic thresholding functions, and the corresponding TV functions. The following results apply to both versions.
@@ 3.2 Theoretical Analysis
We will next demonstrate that ${\mathcal{S}}_{\tau}({\bm{z}})$ provides a close approximation of the proximal operator of TV. First, we show that it is related to the following smoothed versions of $h$. In the anisotropic case, define
$\widetilde{h}^{a}({\bm{z}})=\sum_{i=1}^{nd}\phi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{i}),$ (9)
where $\phi_{\theta}:\mathbb{R}\rightarrow\mathbb{R}$ is defined as
$\phi_{\theta}(y)=\begin{cases}\frac{y^{2}}{2\theta}&\text{for }|y|\leq\theta\\ |y|-\frac{\theta}{2}&\text{for }|y|>\theta\end{cases}.$ (10)
For the isotropic case, define
$\widetilde{h}^{i}({\bm{z}})=\sum_{i=1}^{n}\psi_{\tau 4d}([{\bm{D}}{\bm{z}}]_{G_{i}}),$ (11)
where for ${\bm{y}}\in\mathbb{R}^{d}$ we define
$\psi_{\theta}({\bm{y}})=\begin{cases}\frac{\|{\bm{y}}\|^{2}}{2\theta}&\text{for }\|{\bm{y}}\|\leq\theta\\ \|{\bm{y}}\|-\frac{\theta}{2}&\text{for }\|{\bm{y}}\|>\theta\end{cases}.$ (12)
Note that each $G_{i}$ is the collection of $d$ indices of the vector ${\bm{D}}{\bm{z}}$ corresponding to the finite difference vector at position $i$ of $\bm{z}$}. Related approximations have previously appeared in the TV literature [30, 56, 57, 58, 59, 60], based on the Huber function [61] or Moreau-Yoshida regularization [62, 63].
Proposition 2.
Let $\widetilde{h}^{a}({\bm{z}})$ and $\widetilde{h}^{i}({\bm{z}})$ denote the Huber variants of anisotropic TV and isotropic TV given in Eq. (9) and Eq. (11), respectively. Then,
${\mathcal{S}}^{a}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{a}({\bm{z}})\quad\text{and}\quad{\mathcal{S}}^{i}_{\tau}({\bm{z}})={\bm{z}}-\tau\nabla\widetilde{h}^{i}({\bm{z}}).$
Proposition 2 states that the operator ${\mathcal{S}}_{\tau}({\bm{z}})$ is equivalent to a gradient step with step-size $\tau$ on a smoothed version of TV. A recent work [64] demonstrates that running a single iteration of the dual forward backward algorithm to solve the TV proximal within a FPG algorithm results in an algorithm that converges to a solution of the TV minimized problem. Proposition 2 shows that ${\mathcal{S}}_{\tau}({\bm{z}})$ is related; however, instead of running one iteration of the dual-FB algorithm, ${\mathcal{S}}_{\tau}({\bm{z}})$ corresponds to one iteration of gradient descent on a smoothed version of TV. To decrease Eq. (9) or Eq. (11), the step-size must be at most $1/L$, where $L$ is the Lipschitz constant of the function’s gradient [65] (Section 9.3). For both Eq. (9) and Eq. (11), the Lipschitz constant is $1/{\tau}$, as verified in the Appendix, satisfying the condition.
To further describe the operator as an approximation of the true proximal operator for h{h}, we introduce the concept of the subdifferential [62, 66].
@@
Proposition 3 relies on notions of approximate proximal operators, commonly used in the literature. Specifically, Proposition3 (b) aligns with a “type 3 approximation” in the proximal operator with $\tau\epsilon_{1}$-precision, while Proposition 3 (c) corresponds to a “type 2 approximation” with $\tau\sqrt{2\epsilon_{2}}$-precision [49].
The first part of Proposition 3 demonstrates that the operator functions as a perturbed proximal operator of TV. Specifically, the output of the operator is equivalent to applying the true proximal operator of TV to a perturbed input vector $\bm{z}$}, modified by a vector $\bm{\delta}$}. Notably, the norm of the perturbation $\|{\bm{\delta}}\|_{2}$ is bounded and can be controlled by adjusting $\tau$. From the optimality condition for the proximal operator of true TV regularizer ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$, there is the following well known relationship
${\bm{z}}-{\mathsf{prox}}_{\tau{h}}({\bm{z}})\in\tau\partial\hskip 1.49994pt{h}({\mathsf{prox}}_{\tau{h}}({\bm{z}}))\;.$
The second part of Proposition 3 establishes an analogous connection for the closed-form approximate operator, characterized using the $\epsilon$-subdifferential from Definition 2, where $\epsilon$ is controlled by the proximal scaling parameter $\tau$.
Overall, Proposition 3 implies that the $\widehat{h}$ from Proposition 1 can be seen as an approximation of the true TV function h{h} with controllable accuracy. Notably, by reducing $\tau$, one can achieve an arbitrarily close approximation to the true TV proximal operator. Therefore, our theoretical result supports the use of ${\mathcal{S}}_{\tau}({\bm{z}})$ as an approximation of ${\mathsf{prox}}_{\tau{h}}({\bm{z}})$. As the accuracy of the operator is controlled by $\tau$, it might initially appear that it would only be effective for problems with a very small regularization parameter $\lambda$ for TV. However, this is not the case when ${\mathcal{S}}_{\tau}({\bm{z}})$ is used within iterative algorithms such as APGM and ADMM. As shown in Algorithms 1 and 2, the shrinkage parameter for ${\mathcal{S}}_{\tau}({\bm{z}})$ is set by $\tau=\gamma\lambda$. Thus, one can enhance accuracy by adjusting $\gamma$ rather than altering the TV regularization parameter $\lambda$ . For APGM, $\gamma$ represents the step size, while for ADMM, $\gamma$ serves as the penalty parameter in the augmented Lagrangian.
@@
Table 1: The relative cost, PSNR with respect to the exact TV solution and groundtruth, and acceleration achieved by using ${\mathcal{S}}_{\tau}({\bm{z}})$ (measured as the ratio of the number of iterations when using ${\mathcal{S}}_{\tau}({\bm{z}})$ to total number of sub-iterations when using the exact TV proximal) for various regularization ($\lambda$) and step-size/penalty ($\gamma$) parameters for image denoising and CT reconstruction. For the APGM and ADMM algorithms, $\tau=\gamma\lambda$, where $\gamma$ determines the accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$. The metrics have been averaged across the 10 images in the dataset. Note that $L$ is the Lipschitz constant of $\nabla g$.
APGM Image Denoising
$\lambda=0.25$ $\lambda=0.5$ $\lambda=1$
Step Size ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1e-1 1.751e-02 31.97 16.44 $2.32\times$ 1.214e-01 23.14 17.57 $9.58\times$ 3.875e-01 19.65 17.38 $19.86\times$
1e-2 1.364e-03 48.02 16.72 $0.41\times$ 1.281e-02 34.52 18.57 $1.61\times$ 5.467e-02 25.82 15.68 $2.67\times$
1e-3 1.152e-04 65.90 16.72 $0.09\times$ 1.157e-03 49.16 18.75 $0.38\times$ 5.945e-03 36.31 15.96 $0.67\times$
APGM CT Reconstruction
$\lambda=2.5$ $\lambda=5$ $\lambda=10$
Step Size ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1/(1LL) 9.473e-04 57.28 21.07 $49.50\times$ 2.069e-03 52.08 20.86 $49.21\times$ 4.588e-03 46.48 20.41 $49.71\times$
1/(2LL) 4.609e-04 59.46 21.06 $39.99\times$ 1.002e-03 56.87 20.88 $40.20\times$ 2.208e-03 51.84 20.48 $39.21\times$
1/(4LL) 2.264e-04 62.14 21.07 $33.51\times$ 4.874e-04 60.42 20.88 $32.99\times$ 1.068e-03 56.79 20.51 $31.09\times$
ADMM CT Reconstruction
$\lambda=2.5$ $\lambda=5$ $\lambda=10$
Pen. Par. ($\gamma$) Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel Rel. Err. PSNR (TV) PSNR (GT) Accel
1e-2 3.545e-01 18.49 15.35 $260.90\times$ 4.341e-01 18.10 15.13 $157.57\times$ 4.680e-01 18.28 15.13 $89.96\times$
1e-3 3.767e-02 31.63 19.82 $62.75\times$ 8.291e-02 26.42 18.59 $69.34\times$ 1.788e-01 21.91 16.82 $69.53\times$
1e-4 4.850e-03 40.42 20.59 $11.43\times$ 9.284e-03 38.55 20.37 $10.24\times$ 1.989e-02 35.13 19.82 $9.90\times$
@@ 4 Numerical Validation
In this section, we numerically validate Proposition 3 by comparing the reconstruction performance of iterative proximal algorithms when ${\mathcal{S}}_{\tau}({\bm{z}})$ replaces the true proximal operator for isotropic total variation. For experimental validation of the efficiency and timing of these operators, we refer the reader to [8, 9, 10]. Specifically, we run APGM and ADMM on a simulated low-angle computed tomography (CT) image reconstruction problem. The APGM and ADMM algorithms using ${\mathcal{S}}_{\tau}({\bm{z}})$ instead of the true proximal operator are shown in Algorithms 1 and 2, respectively. The accuracy of ${\mathcal{S}}_{\tau}({\bm{z}})$ can be directly controlled by $\gamma$ since $\tau=\gamma\lambda$, where $\gamma$ is the step size and penalty parameter, for APGM and ADMM algorithms, respectively. To calculate the true proximal operator, we use the iterative Fast Projected Gradient (FPG) algorithm [5]. Additionally, we investigate the performance of the closed-form approximate operator relative to true proximal operator of TV for an image denoising task. The numerical evaluations are performed using 10 piecewise-smooth foam phantoms generated by the XDesign Python package [68] and the implementationsSee classes AnisotropicTVNorm and IsotropicTVNorm. of ${\mathcal{S}}_{\tau}({\bm{z}})$ in the SCICO Python package [69]. All algorithms are run until the stopping criteria